# Exploring Young Stellar Variability with Gaia DR3 light curves

### v1

@@ 5.2 Amplitude and timescales
We further complemented our results analysis with peak-to-peak (ptp) amplitudes and timescales measurements. We used the min-max amplitude as a measure of the ptp amplitude rather than the 5-95 percentile amplitude more commonly employed in the literature, as we found the latter to underestimate ptp amplitudes when applied to GDR3 sparse sampling. Throughout the discussion of results, we refer to timescales in the range d as rotation timescales. Timescales larger than d and less than about half of the duration ( days) of the observations are termed intermediate timescales. Timescales longer than half of the duration of the light curve are referred to as long-trend timescales.
Fig. 11 shows the distribution of amplitudes as a function of timescales for our sample, revealing a bimodal distribution of timescales with a population of sources with rotation-timescales () and a second population of variables with intermediate timescales ( of the sample).
@@
As detailed in Sect. 4.2.2 and App. B, a combination of factors such as the sparse sampling of GDR3 and the SF discretisation process makes the SF method inefficient for detecting aperiodic timescales in the range d. Hence, if the GLS method was inefficient at detecting long timescales, this bi-modality could be explained by the detection limits introduced by the combined use of these two methods. Conversely, Fig. 11 shows that the distribution of timescales measured with the GLS method is itself bi-modal, with 28% of GLS-derived values at intermediate-timescales. Additionally, albeit much reduced in numbers, the SF timescale distribution accounts for of rotation timescales. Another possible explanation would be that the intermediate-timescale population is the result of the spurious variability discussed in Sect.3.3. However, we have addressed and mitigated the occurrence of residual spurious variability in our sample in App. 29 and Sect. 4.3, where we discarded sources with timescales in the ranges typically susceptible to spurious variability. Finally, when comparing our results with previous studies using ZTF light curves (Sect. 6.3.2), we found that those previous studies also find a hint to similar bi-modality in timescales, with vs incidence of rotation and intermediate timescales. We are therefore confident that the observed bi-modality is real.
The two populations also show distinctive behaviour in ptp amplitudes, with rotation timescales showing average amplitudes of 0.18 mag (standard deviation 0.26) and intermediate timescales at 0.29 mag (0.38), and these differences are also supported by a Kolmogorov–Szmirnov (KS) test.
5.3 Asymmetry and Periodicity metrics in the context of YSO’s observed properties
@@
5.3.1 Accreting vs. non-accreting sources
We used the data described in Sect. 2.2.2 to classify sources as Classical T Tauri Stars (CTTS) based on their H emission and as actively accreting sources. For equivalent width measurements, we employed spectral type-dependent criteria by Barrado y Navascués & Martín (2003) converted into Gaia - colours using the Pecaut & Mamajek (2013) empirical sequence. For full widths at 10% of the peak flux measurements, we used the White & Basri (2003) criterion to distinguish between CTTS and weak-lined TTS (WTTS). For 64 sources, we used qualitative flags that report H in emission (from Pettersson et al., 2014). When multiple H measurements were available, we considered that one measurement indicating a CTTS classification was enough to grant the source a classification as such. Specifically for EW(H) measurements from ESP-ELS Gaia DR3 we have restricted our analysis to 7,318 sources with the absolute value of reported EW(H) larger than at least 5 times reported uncertainties. This decision was motivated by the limitations due to / spectral low resolution and its representation as Hermite function bases (; De Angeli et al., 2023), where the identification of sources with large values of EW(H) are reliable, but where low accreting and non-accreting sources with EW(H) values closer to the stellar continuum may be confused within uncertainties.
@@
In total, we could assign CTTS/WTTS labels to 13,802 sources in the YSO sample with at least one H measurement, 7 738 in the variable sample, and 2 829 in our primary variable sample with both Q&M derived, with 1 107 CTTSs and 1 722 WTTSs. The Q vs. M distribution for these sources is shown in Fig. 12, where we further distinguish between disk-bearing and diskless WTTS based on their -index (see Sect. 5.3.2). The sample of sources with H information has a lower incidence of aperiodic sources, with only of these compared to in the full variable sample.
@@
The Q&M distribution for diskless WTTS peaks at Q and M, with the distribution in Q shifted toward the intersection between strictly and quasiperiodic sources and a narrower distribution of M values concentrated toward symmetric M indexes. The subsample of actively accreting sources (CTTS) is mainly quasiperiodic, with their distribution peaking at Q and M. Compared to the diskless WTTS distribution, the CTTS is shifted towards larger Q values, with a much wider distribution of M values including more significant fractions of dippers and bursters. For disk-bearing WTTSs, Q and M, with the distribution presenting intermediary properties between the former two groups. Disk-bearing WTTSs show a smaller relative fraction of strictly periodic sources compared to diskless WTTS as well as an excess of quasiperiodic dippers with larger M-values compared to CTTSs.
@@
To further support results discussion, Table 3 includes the CTTS frequency. This is defined as the percentage of CTTS sources among sources with H measurements and is interpreted here as a measure of the fraction of actively accreting sources within a sample. However, we note that H measurements are only possible for optically visible sources, thus this proxy will inevitably miss deeply embedded sources. Furthermore, we also note that the accretion process and the resulting H-emission are also highly variable (e.g. Sousa et al., 2016). Hence, although strong H emission indicates ongoing accretion, the opposite is not always true, as the source may still be sporadically accreting over longer timescales although not in the specific season observed. CTTS fraction must thus be interpreted as a bottom limit of the true fraction of actively accreting sources.
5.3.2 Circumstellar disks
@@
To gauge the influence of circumstellar disks on our variability results, we also examined the indices for our sources. The index is a measure of the slope of the IR spectral energy distribution (SED) and quantifies the contribution to the SED of infrared radiation re-emitted by material around YSOs. To estimate we follow a similar procedure as the one in the NEMESIS YSO catalogue: We used the Vizier Photometry viewer tool to retrieve all previously published photometric data available at the Centre de données astronomiques de Strasbourg (Genova et al., 2000; Ochsenbein et al., 2000), and used data in the wavelength range m to fit a linear model to measure . Next, we employ an adaptation of the standard classification scheme for YSOs based on their (Lada, 1987; Andre et al., 1993; Greene et al., 1994), where, to facilitate our discussion, we name subclasses based on the dominant contributor to their SED: sources with significant contribution of their envelope have (483 sources in the primary sample), sources with a thick disk have (3 856 sources), sources with thin disks have (1 741 sources) and diskless sources have , (2 320 sources). Fig. 13 shows the Q vs. M distributions for each of these samples. Table 3 includes the frequency of sources with -based evidence for the existence of an inner disk, which we define as the total number of sources classified as Envelope or Thick disk, divided by the total number of sources with derived -index.
@@
KS tests between envelope and thick disk samples suggest that these samples share the same mother Q&M distribution. However, when examined in terms of the ratio between rotation and intermediate timescales, we find a 0.7 ratio for the envelope sample compared to a 1.0 ratio for the thick disk sample. Generally, Fig. 13 suggests a gradual transition towards a distribution of M-indices that favours symmetry and a shift toward enhanced periodicity for more evolved YSO populations.
@@
Several of the variability mechanisms in YSOs stem from the star-disk interaction process, where the young star magnetically interacts with its inner disk. Hence a relevant quantity included in Table 3 is the percentage of sources with index indicative of the presence of a inner disk, which we define as and includes both sources with a thick disk or an envelope. We stress that these quantity also represent the bottom limit of the true frequencies. Our are derived based on heterogeneous data available from previous publications and thus may be biased toward a lack of mid- and far-infrared data, which are only available all-sky within the context of WISE (Wright et al., 2010), with limited coverage for .
@@ 6 Discussion
The effectiveness of the Q&M-metrics for characterising YSO variability has been previously demonstrated in the literature with successful applications to high-cadence data from space telescopes and data with longer timespan but from ground-based telescopes. With the present study, we contributed to this body of research by demonstrating the applicability of the Q&M-metrics to Gaia’s sparse, but high-precision, long-term data. With this applicability demonstrated, in this section we proceed to discuss the interpretation of our results given the morphological classes defined with Q&M.
@@ 6.1 Incidence of morphological classes across surveys
In this section, we discuss the incidence of morphological classes in different surveys. Here and in Sect. 6.2, literature comparison is discussed at the population level. A comparison of some individual sources in common between the surveys is presented in Sect. 6.3. Table 4 shows a comparison of the frequencies of variability morphology in different studies. Morphological classes from the literature are taken as reported, without review of their Q&M thresholds. Fig. 14 shows the distribution of amplitudes versus timescales for each morphological class. This includes our primary and secondary samples, as well as the main samples from the literature combined by type of survey, including high-cadence surveys (CoRoT and K2 Cody et al., 2014; Cody & Hillenbrand, 2018; Venuti et al., 2021; Cody et al., 2022) and ground-based results (from ZTF; Hillenbrand et al., 2022; Wang et al., 2023; Jiang & Hillenbrand, 2024). Detailed discussion of the properties of each morphological class are later presented in Sect. 6.2.
This Work | Previous Studies | ||
Class | Full Sample | High-cadence | ZTF |
P | 1.0 | 6.7 | 14.5 |
QPS | 33.7 | 28.3 | 35.6 |
QPD | 19.3 | 16.1 | 13.8 |
APD | 3.9 | 15.7 | 10.0 |
B | 18.6 | 17.6 | 8.9 |
S | 5.3 | 12.8 | 14.0 |
EB | 0.5 | 0.2 | - |
L | 17.6 | 2.6 | 3.2 |
@@
Discussion of results must thus be carried out while acknowledging survey-to-survey differences. First, in terms of sample size: our Gaia study included 9 824 YSOs in the primary sample and 18,355 in the secondary sample, while ZTF and high-cadence surveys included 1 031 and 549 sources, respectively. With much smaller sample sizes, previous studies heavily relied on visual inspection of light curves for classification, with Q&M thresholds adjusted by eye to maximise its correspondence with visually attributed morphological classes. Although infeasible in large-scale surveys such as Gaia, this empirical approach adopted in previous studies suggests that survey-to-survey adjustments of Q&M thresholds can help account for surveys’ sampling limitations and may improve purity within morphological classes. In Sect. 4.4, we adopted Cody et al. (2014) original Q&M thresholds for morphological classification. As further discussed in Sect. 6.2, our results suggest that these thresholds led to some contamination of symmetric light curves into asymmetric classes within our sample. With the ground truth of visual inspection further hindered by Gaia’s sparse light-curves, an alternative approach would be to review Q&M thresholds based on simulated light curves. Although beyond the scope of the present study, our team has been exploring this approach as part of our preparations for Gaia DR4.
@@
There are also source selection differences, where high-cadence surveys examined only disk-bearing sources, while both our results and ZTF’s include all types of YSOs. Moreover, previous studies focused on individual star-forming regions, with much narrower age-distributions in comparison with the GDR3 all-sky sample. Notably, Cody et al. (2022) attempted to interpret the frequency distribution of morphological classes in different star-forming regions in light of their age. In our case, this type of interpretation is hindered by the lack of previous constraints in the age distribution of the GDR3 YSO catalogue. However, our investigations of -indices in Sect. 5.3.2 revealed that 42% and of our sample of variables had no indication of the presence of a disk or only a thin-disk contribution to their SEDs.
@@
Given the enhanced fraction of diskless YSOs in our sample, cool spot variability, typically associated with P and QPS sources, is expected to be the most common in our sample. Although QPS is in fact the most common variability class, Table 4 shows that the frequency of P-variables is much lower than in previous studies. As further discussed in Sect. 6.2.1, beyond cadence and underlying noise, this can be attributed to cool-spot variability being unstable at timescales of the order of GDR3 observations. We note that cool-spot variability is a central feature of ”solar-like variables”, which are themselves a subgroup of Gaia DR3 variability classes (Distefano et al., 2022), where the authors report that between 53 to 72% of their 474 026 variables show unstable spot-variability (amplitude variation) during the observed time covered by GDR3. This helps explain not only our under-detection of the P-variable compared to previous studies, but also contributes to our enhanced detection of QPS variables. Table 4 also indicates that ZTF studies, which likewise include diskless sources, show a much larger sample of P and QPS sources in comparison to high-cadence surveys. However, ZTF studies use much larger Q thresholds to account for the higher noise from ground-based observations.
@@
The solar-like variables identified by Distefano et al. (2022) focus on late-type stars, which include some relatively young sources, while being disjoint from the GDR3 YSO sample. This suggests that a certain fraction of young P and QPS variables may be lost to their sample. To quantify sources lost to GDR3’s sample of solar-like variables, we further examined the NEMESIS YSO catalogue for the Orion Star Formation Complex (Roquette et al., 2025) which contains 27 879 YSO candidates. The NEMESIS YSO sample included 5 231 sources in common with Gaia’s YSO sample, of which 111 were classified as P and 2 131 as QPS, and further 152 sources in common with Gaia’s solar-like variables, suggesting that we miss about of cool-spot variables. Similarly, EBs are also a subclass of variables in Gaia, with Mowlavi et al. (2023) reporting more than 2 million EB candidates as part of DR3. A similar comparison to the NEMESIS catalogue suggests that our EB-detection rate might be reduced by half.

### v2

@@ 5.2 Amplitude and timescales
We further complemented our results analysis with peak-to-peak (ptp) amplitudes and timescales measurements. We used the min-max amplitude as a measure of the ptp amplitude rather than the 5-95 percentile amplitude more commonly employed in the literature, as we found the latter to underestimate ptp amplitudes when applied to GDR3 sparse sampling. Throughout the discussion of results, we refer to timescales in the range d as “rotation timescales”. Timescales larger than d and less than about half of the duration ( days) of the observations are termed “intermediate timescales”. Timescales longer than half of the duration of the light curve are referred to as “long-trend timescales”. Fig. 11 shows the distribution of amplitudes as a function of timescales for our sample, revealing a bimodal distribution of timescales with a population of sources with rotation-timescales () and a second population of variables with intermediate timescales ( of the sample).
| FrequencyaaaaValues in parentheses are standard error estimates for the disc frequencies. | Spectral TypebbbbFrequency of spectral types as a function of variability class | |||||||||
| Class | count | thick disc | general disc | CTTS | ts-ratio | ptp | M | K | G | AB-F |
| Full sample | ||||||||||
| P | 286 | 14.3 (2.1) | 50.0 (3.0) | 13.7 (2.8) | 13.4 | 0.19 | 77.3 (2.5) | 12.9 (2.0) | 1.7 (0.8) | 8.0 (1.6) |
| QPS | 9514 | 20.0 (0.4) | 57.6 (0.5) | 14.8 (0.6) | 2.6 | 0.15 | 71.2 (0.5) | 15.3 (0.4) | 4.0 (0.2) | 9.5 (0.4) |
| QPD | 5439 | 39.1 (0.7) | 67.3 (0.6) | 24.7 (1.1) | 1.8 | 0.35 | 65.4 (0.6) | 18.6 (0.5) | 5.4 (0.3) | 10.6 (0.4) |
| APD | 1103 | 56.6 (1.5) | 79.9 (1.2) | 39.7 (3.7) | 0.4 | 0.53 | 66.2 (1.4) | 19.7 (1.2) | 5.7 (0.7) | 8.4 (0.8) |
| B | 5252 | 21.0 (0.6) | 56.7 (0.7) | 22.0 (1.1) | 2.0 | 0.16 | 59.8 (0.7) | 24.3 (0.6) | 3.2 (0.2) | 12.7 (0.5) |
| S | 1490 | 27.9 (1.2) | 65.9 (1.2) | 23.6 (2.6) | 0.8 | 0.19 | 71.4 (1.2) | 17.8 (1.0) | 3.2 (0.5) | 7.6 (0.7) |
| EB | 129 | 15.5 (3.2) | 56.6 (4.4) | 11.4 (3.6) | 23.6 | 0.20 | 85.3 (3.1) | 9.3 (2.6) | 0.8 (0.8) | 4.6 (1.9) |
| L | 4966 | 21.1 (0.6) | 48.5 (0.7) | 11.0 (1.1) | - | 0.16 | 44.3 (0.7) | 22.9 (0.6) | 7.7 (0.4) | 25.1 (0.6) |
| Primary sample | ||||||||||
| P | 199 | 17.1 (2.7) | 46.2 (3.5) | 14.6 (3.1) | 26.6 | 0.22 | 81.4 (2.8) | 8.6 (2.0) | 2.0 (1.0) | 8.0 (1.9) |
| QPS | 2362 | 41.9 (1.0) | 65.5 (1.0) | 36.6 (1.6) | 1.4 | 0.31 | 60.5 (1.0) | 22.2 (0.9) | 7.1 (0.5) | 10.2 (0.6) |
| QPD | 2886 | 61.7 (0.9) | 81.4 (0.7) | 40.6 (1.6) | 1.3 | 0.57 | 65.5 (0.9) | 20.5 (0.8) | 6.1 (0.4) | 7.9 (0.5) |
| APD | 822 | 68.0 (1.6) | 85.2 (1.2) | 48.2 (4.2) | 0.4 | 0.67 | 63.4 (1.7) | 21.9 (1.5) | 7.0 (0.9) | 7.7 (0.9) |
| B | 1505 | 44.8(1.3) | 69.4 (1.2) | 45.6 (2.2) | 1.3 | 0.31 | 68.4 (1.2) | 18.2 (1.0) | 4.3 (0.5) | 9.1 (0.7) |
| S | 601 | 42.8 (2.0) | 70.9 (1.9) | 49.0 (4.9) | 0.4 | 0.30 | 62.1 (2.0) | 23.3 (1.7) | 4.8 (0.9) | 9.8 (1.2) |
| EB | 95 | 15.8 (3.7) | 51.6 (5.1) | 14.1 (4.3) | 91.0 | 0.24 | 84.2 (3.7) | 8.4 (2.8) | 1.1 (1.1) | 6.3 (2.5) |
| L | 1354 | 40.0(1.3) | 63.1 (1.3) | 33.6 (3.2) | 0.37 | 48.6 (1.4) | 26.3 (1.2) | 8.2 (0.7) | 16.9 (1.0) | |
| Secondary sample | ||||||||||
| P | 87 | 8.0 (2.9) | 58.6 (5.3) | 8.7 (5.9) | 5.8 | 0.10 | 67.8 (5.0) | 23.0 (4.5) | 1.1 (1.1) | 8.0 (2.9) |
| QPS | 7152 | 12.8 (0.4) | 55.0 (0.6) | 6.1 (0.5) | 3.3 | 0.09 | 74.7 (0.5) | 13.1 (0.4) | 3.0 (0.2) | 9.2 (0.3) |
| QPD | 2553 | 13.4 (0.7) | 51.5 (1.0) | 4.5 (0.8) | 2.5 | 0.09 | 65.2 (0.9) | 16.5 (0.7) | 4.6 (0.4) | 13.7 (0.7) |
| APD | 281 | 23.1 (2.5) | 64.4 (2.9) | 3.0 (3.0) | 0.7 | 0.13 | 74.4 (2.6) | 13.5 (2.0) | 1.8 (0.8) | 10.3 (1.8) |
| B | 3747 | 11.4 (0.5) | 51.6 (0.8) | 7.0 (0.9) | 2.4 | 0.10 | 56.4 (0.8) | 26.8 (0.7) | 2.6 (0.3) | 14.2 (0.6) |
| S | 889 | 17.8 (1.3) | 62.5 (1.6) | 7.4 (2.0) | 1.1 | 0.12 | 77.8 (1.4) | 14.0 (1.2) | 2.1 (0.5) | 6.1 (0.8) |
| EB | 34 | 14.7 (6.1) | 70.6 (7.8) | - | 6.8 | 0.10 | 88.2 (5.5) | 11.8 (5.5) | - | - |
| L | 3612 | 14.0 (0.6) | 43.0 (0.8) | 2.5 (0.7) | - | 0.09 | 42.6 (0.8) | 21.6 (0.7) | 7.6 (0.4) | 28.2 (0.7) |
@@
As detailed in Sect. 4.2.2 and App. B, a combination of factors such as the sparse sampling of GDR3 and the SF discretisation process makes the SF method inefficient for detecting aperiodic timescales in the range d. Hence, if the GLS method was inefficient at detecting long timescales, this bi-modality could be explained by the detection limits introduced by the combined use of these two methods. Conversely, Fig. 11 shows that the distribution of timescales measured with the GLS method is itself bi-modal, with 28% of GLS-derived values at intermediate-timescales. Additionally, albeit much reduced in numbers, the SF timescale distribution accounts for of rotation timescales. Another possible explanation would be that the intermediate-timescale population is the result of the spurious variability discussed in Sect.3.3. However, we have addressed and mitigated the occurrence of residual spurious variability in our sample in App. 31 and Sect. 4.3, where we discarded sources with timescales in the ranges typically susceptible to spurious variability. Finally, when comparing our results with previous studies using ZTF light curves (Sect. 6.3.2), we found that those previous studies also find a hint to similar bi-modality in timescales, with vs incidence of rotation and intermediate timescales. We are therefore confident that the observed bi-modality is real. The two populations also show distinctive behaviour in ptp amplitudes, with rotation timescales showing average amplitudes of 0.18 mag (standard deviation 0.26) and intermediate timescales at 0.29 mag (0.38), and these differences are also supported by a KS test.
5.3 Asymmetry and periodicity metrics in the context of the observed YSO properties
@@
5.3.1 Accreting versus non-accreting sources
@@
In App. C.1, we classified sources as actively accreting Classical T Tauri Stars (CTTS) or non-accreating weak-lined TTS (WTTS) based on their H emission. In total, we could assign CTTS/WTTS labels to 13 802 GDR3 YSOs, with 7 738 in the variable sample. Among 2 829 sources in the primary sample with both Q and M derived, 1 107 were CTTSs and 1 722 were WTTSs. Their Q-M distribution is shown in Fig. 12, where we further distinguish between disc-bearing and discless WTTS based on their index (Sect. 5.3.2). Sources with H information have a lower incidence of aperiodic variability () compared to the full variable sample ().
@@
The Q-M distribution for discless WTTS peaks at and , with Q values shifted toward the intersection between strictly and quasi-periodic sources, and a narrower distribution of M values concentrated toward symmetric M indices. The subsample of actively accreting sources (CTTS) is mainly quasi-periodic, with their distribution peaking at and . Compared to the discless WTTS distribution, CTTSs are shifted towards larger Q values, with a much wider distribution of M values including more significant fractions of dippers and bursters. For disc-bearing WTTSs, and , with the distribution presenting intermediary properties between the former two groups. Disc-bearing WTTSs show a smaller relative fraction of P sources compared to discless WTTS as well as an excess of QPDs with higher M values than CTTSs.
@@
To further support the discussion of the results, Table 1 includes the CTTS frequency defined as the percentage of CTTS sources among all sources with H measurements, and interpreted here as a measure of the fraction of sources actively accreting within a sample. We note that H measurements are limited to optically visible sources. Consequently, this proxy will inevitably miss deeply embedded sources. Furthermore, we also note that the accretion process and the resulting H-emission are also highly variable (e.g. Sousa2016A&A...586A..47S). Hence, although strong H emission indicates ongoing accretion, the opposite is not always true, as sources may still be sporadically accreting over longer timescales although not in the specific season observed. CTTS fraction must thus be interpreted as a bottom limit of the true fraction of actively accreting sources.
5.3.2 Circumstellar discs
@@
To gauge the influence of circumstellar discs on our results, we examined the indices for our sources. The index is a measure of the slope of the IR spectral energy distribution (SED) and quantifies the contribution to the SED of infrared radiation re-emitted by material around YSOs. In App. C.2, we were able to derive for 27 653 sources (98% of our final sample). To facilitate our discussion, we used to classify sources according to the dominant contributor to their SED: 712 sources in the primary sample and 697 in the secondary had a significant envelope contribution; 4 296 (primary) and 2 131 (secondary) sources had a thick disc; 2 021 and 6 722 sources had a thin discs; and 2 795 and 8 836 sources were likely discless. Fig. 13 shows the Q-M distributions for each of these samples. Table 1 includes the frequency of sources with -based evidence for the existence of an inner disc, which we define as the total number of sources classified as envelope or thick disc, divided by the total number of sources with index.
@@
KS tests between envelope and thick disc samples suggest that these samples may share the same mother Q-M distribution. However, when examined in terms of the ratio between rotation and intermediate timescales, we find a 0.7 ratio for the envelope sample compared to a 1.0 ratio for the thick disc sample. Generally, Fig. 13 suggests a gradual transition towards a distribution of M indices that favours symmetry and a shift toward enhanced periodicity for more evolved YSO populations.
@@
Several of the variability mechanisms in YSOs stem from the star-disc interaction process, where the young star magnetically interacts with its inner disc. Hence a relevant quantity included in Table 1 is the percentage of sources with index indicative of the presence of a inner disc, which we define as and includes both sources with a thick disc or an envelope. We also report the fraction of sources with evidence for circumstellar material (namely general disc-fraction), which we define as the fraction of sources or that had EW(H) indicative of accretion (Sect. 5.3.1.). Our are derived based on heterogeneous data available from previous publications and thus may be biased toward a lack of mid- and far-infrared data, which are only available all-sky within the context of WISE (Wright2010_WISE), with of our compiled SEDs including only data below . We therefore stress that reported disc frequencies represent a lower bound on the actual frequencies. For example, the thick disc fraction for the full sample goes from 25.5% to 37.3% when considering only sources with SEDs populated beyond . For a discussion of how using SEDs without data beyond affects studies based on , see roquette2024, who report a 25.2% impurity among thin disc sources identified from such undersampled SEDs.
5.3.3 Spectral types
Fig. 15 shows the distribution of spectral types for our sample, estimated in App. C.3, illustrating that of our sample is composed of K and M sources, with a smaller incidence () of intermediate mass YSOs. In Table 1 we present the incidence of spectral types in each variability class, and in Fig. 15, the incidence of variability classes for each spectral type. The latter, in particular, illustrates two trends as a function of spectral type. Earlier spectral types are more frequently classified as L variables, whereas QPS sources occur more often at later spectral types. We also note that B, QPD and APD also vary with spectral type, although variations are much smaller and show a less clear trend. In our discussion of the results, we group A, B and F type objects as intermediate-mass or early-type YSOs. G , K, and M type objects are treated as late-type sources, with K and M type sources distinguished when relevant.
@@ 6 Discussion
The effectiveness of the Q-M metrics for characterising YSO variability has already been demonstrated in the literature with successful applications to both high-cadence space-based and longer-term ground-based observations. In this study, we extend prior work by showing that Q–M metrics can be applied to Gaia’s sparse yet highly precise long-term data. Having established this, we next consider how to interpret our findings within the Q–M morphological classification scheme.
@@ 6.1 Incidence of morphological classes across surveys
Table 2 compares the population-level frequencies of variability morphologies reported in different works. Morphological classes from the literature are taken as reported, without review of their Q-M thresholds. Some previous studies included extra classes for multi-periodic and non-variables, which we did not consider here and we estimated their class incidences excluding these two categories. We thus stress that these frequencies reflect the incidence of morphological classes among samples of variable stars. Fig. 16 presents amplitudes versus timescales distributions for each morphological class. This includes our primary and secondary samples, as well as the main samples from the literature combined by type of survey, including high-cadence surveys (CoRoT and K2; Cody_14; Cody_2018; Venuti_2021; Cody_2022) and ground-based results (ZTF; Hillenbrand2022; Wang2023; JiangHillenbrand2024). Detailed discussion of each morphological class are later presented in Sect. 6.2.
| This Work | Previous StudiesaaaaThe reported incidences from previous studies refer to the primary variability classifications and were recalculated to account for the total number of variables in P, QPS, QPD, QPD, B, S, EB, and L. | ||||
| Full | Pri- | Sec- | High- | ||
| Class | Sample | mary | ondary | cadence | ZTF |
| P | 1.0 | 2.0 | 0.5 | 6.7 | 14.5 |
| QPS | 33.7 | 24.0 | 39.0 | 28.3 | 35.6 |
| QPD | 19.3 | 29.4 | 13.9 | 16.1 | 13.8 |
| APD | 3.9 | 8.4 | 1.5 | 15.7 | 10.0 |
| B | 18.6 | 15.3 | 20.4 | 17.6 | 8.9 |
| S | 5.3 | 6.1 | 4.8 | 12.8 | 14.0 |
| EB | 0.5 | 1.0 | 0.2 | 0.2 | - |
| L | 17.6 | 13.8 | 19.7 | 2.6 | 3.2bbbbThe L incidence reported for ZTF (3.2%) does not account for sources classified as such in their secondary classifications (). |
@@
Discussion of results must acknowledge survey-to-survey differences. First, in terms of sample size: our study included 9 824 YSOs in the primary and 18,355 in the secondary sample, while ZTF and high-cadence surveys included 1 031 and 549 sources, respectively. With much smaller sample sizes, previous studies relied heavily on visual inspection of light curves for classification, with Q-M thresholds adjusted by eye to maximise correspondence with visually assigned morphological classes. Although infeasible in large-scale surveys such as Gaia, this empirical approach adopted in previous studies suggests that survey-to-survey adjustments of Q-M thresholds can help account for surveys’ sampling limitations and may improve purity within morphological classes. In Sect. 4.4, we adopted Cody_14 original Q-M thresholds for classification. As further discussed in Sect. 6.2, our results suggest that these thresholds led to some contamination of symmetric light curves into asymmetric classes within our sample. With the ground truth of visual inspection further hindered by the sparse light curves of Gaia, an alternative approach would be to review Q-M thresholds based on simulated light curves. Although beyond the scope of the present study, our team has been exploring this approach as part of our preparations for Gaia DR4.
High-cadence studies typically have photometric precision of mag, compared to mag for ZTF, implying that the ZTF sample is biased towards higher-amplitude variables. In contrast, our GDR3 sample has typical photometric uncertainties mag for mag, and mag up to mag. Based solely on photometric precision, our primary sample should most closely match ZTF, while for mag, the full sample should best resemble high-cadence surveys. The incidence of morphological classes for the former is shown in Table 2, the latter is omitted, as the results were roughly equivalent to the full sample. However, as discussed in this section, we speculate that cadence and sample selection play a stronger role in explaining different results at the population level across surveys.
@@
There are also source selection differences, where high-cadence surveys examined only disc-bearing sources, while ours and ZTF results include all YSO types. Moreover, previous studies focused on individual star-forming regions, with much narrower age distributions in comparison with the GDR3 all-sky sample. Notably, Cody_2022 attempted to interpret the frequency distribution of morphological classes in different star-forming regions in light of their age. In our case, this type of interpretation is hindered by the lack of previous constraints in the age distribution of the GDR3 YSO catalogue. However, our investigations of -indices in Sect. 5.3.2 revealed that 42% and of our sample of variables had no indication of the presence of a disc or only a thin-disc contribution to their SEDs.
@@
Given the enhanced fraction of discless YSOs in our sample, cool spot variability, typically associated with P and QPS variability, is expected to be the most common in our sample. Although QPS is in fact the most common variability class, Table 2 shows that P frequency is much lower than in previous studies. Beyond cadence and underlying noise, we further discuss in Sect. 6.2.1 that this can be attributed to cool-spot variability being unstable at timescales of the order of GDR3 observations. We note that cool-spot variability is a central feature of “solar-like variables”, which are themselves a subgroup of GDR3 variability classes, with Distefano2022 reporting that 53–72% of their 474 026 variables show unstable spot-variability (amplitude variation) during the observed time covered by GDR3. This helps explain not only our under-detection of P variables compared to previous studies, but also contributes to our enhanced detection of QPS variables. Table 2 also indicates that ZTF studies, which also include discless sources, show a much larger sample of P and QPS sources in comparison to high-cadence surveys. However, ZTF studies use much larger Q thresholds to account for the higher noise from ground-based observations.
@@
The solar-like variables reported by Distefano2022 focus on late-type stars, including some relatively young objects, while being disjoint from the GDR3 YSO sample. This suggests that a certain fraction of young P and QPS variables may be lost to their sample. To quantify sources lost to the GDR3 sample of solar-like variables, we further examined the NEMESIS YSO catalogue for the Orion Star Formation Complex (roquette2024). Of 27 879 YSO candidates, the NEMESIS sample included 5 231 sources in common with the Gaia YSO sample, of which 111 were classified as P and 2 131 as QPS, and a further 152 sources in common with Gaia solar-like variables, suggesting that we miss about of cool-spot variables. Similarly, EBs are also a subclass of variables in Gaia, with Mowlavi2023A&A...674A..16M reporting more than 2 million EB candidates as part of DR3. A similar comparison to the NEMESIS catalogue suggests that our EB detection rate might be reduced by half.